OpenStudy (anonymous):

Find an equation of the tangent line to the graph of f(x) at x=-3 if f(-3)=2 and f'(-3)=5 I am not coming up with the answer that the book gives and need help on the proper steps. Thanks!

OpenStudy (mertsj):

When it tells you that f'(-3)=5 it is giving you the slope of the tangent line. When it tells you that f(-3)=2 it is giving you a point on the tangent. So now you know the slope of the line and a point on the line. Use point-slope form and find the equation of the line.

OpenStudy (anonymous):

@Mertsj I thanks, here is what I did wrong. y-2=5x+3 = y=5x+5 Should have done it as: y-2=5(x+3) =y-2=5x+15 =y=5x+17 Pesky brackets will be my downfall.

OpenStudy (mertsj):

yep